### v1

# Twilight: Adaptive Attention Sparsity with Hierarchical Top- Pruning

Abstract
Leveraging attention sparsity to accelerate long-context large language models (LLMs) has been a hot research topic. However, current algorithms such as sparse attention or key-value (KV) cache compression tend to use a fixed budget, which presents a significant challenge during deployment because it fails to account for the dynamic nature of real-world scenarios, where the optimal balance between accuracy and efficiency can vary greatly. In this paper, we find that borrowing top- sampling (nucleus sampling) to sparse attention can surprisingly achieve adaptive budgeting. Based on this, we propose Twilight, a framework to bring adaptive sparsity to any existing sparse attention algorithm without sacrificing their accuracy. Empirical results show that Twilight can adaptively prune at most 98% of redundant tokens, leading to acceleration in self-attention operations and acceleration in end-to-end per token latency in long context LLM decoding.
1 Introduction
Large language models (LLMs) with long-context capabilities have revolutionized a wide array of natural language processing tasks, such as retrieval-based applications, document summarizationĀ (Bai etĀ al., 2024), and code generationĀ (Jain etĀ al., 2024). The increasing availability of models supporting context windows up to 1M to 10M tokensĀ (Yang etĀ al., 2025; Team etĀ al., 2024) highlights the growing potential of these advancements. For instance, video language models (VLMs)Ā (Wang etĀ al., 2024) that process video often require tens of thousands of tokens. Similarly, large reasoning modelsĀ (DeepSeek-AI etĀ al., 2025; Team etĀ al., 2025), which are rapidly growing in popularity, frequently demand substantial token lengths to enable chain-of-thought (CoT) reasoning. Consequently, the importance of long-context LLMs is increasing rapidly to meet the needs of these sophisticated applications.
Despite the transformative potential of long-context LLMs, they come with substantial computational and memory costsĀ (Zhang etĀ al., 2023; Tang etĀ al., 2024b; Xiao etĀ al., 2024), primarily driven by the attention mechanism. In the decoding stage, the latency is primarily attributed to memory bandwidth limitations, as the KV cache must be repeatedly loaded. This memory-bound nature of attention leads to increased latency since the KV cache sizes grow as the sequence becomes longer. Furthermore, the substantial size of the KV cache significantly increases GPU memory consumption, compounding the challenges of scaling long-context LLMs.
Previous research has extensively investigated the use of attention sparsity (KV cache sparsity) to accelerate long-context inference, both during the prefilling and decoding stages. The core idea is to compute approximate attention on a subset of tokens, often referred to as ācritical tokens" or āheavy hitters"Ā (Zhang etĀ al., 2023). In practical deployments of these algorithms, it is necessary to specify the number of selected tokens, denoted as , commonly referred to as the KV cache budget. A top- operation is required to identify the indices of the critical tokens that correspond to the top- highest estimated scores. As previously mentioned, a smaller significantly reduces the I/O operations, while a larger retains more contextual information, thereby minimizing accuracy loss.
However, identifying the optimal values for where both accuracy and efficiency are optimized is inherently challenging due to two major reasons: (a) Saturation points are runtime dynamic. Previous worksĀ (Wu etĀ al., 2024; Xiao etĀ al., 2024) have demonstrated that some heads, referred to as retrieval heads, are trained to extract important information from long contexts, while others focus only on local information. As shown in FigureĀ 1, the distribution of attention weights may vary across different attention heads. Some attention distributions concentrate on a small subset of tokens, which we refer to as focused attention. In contrast, some attention distributions are flatter, where many tokens have similar attention weights; we define this as diffuse attention. For focused attention, using a fixed token budget for Top- attention often leads to over-selection, as only a few tokens are sufficient to accumulate sufficient attention weights. Similarly, for diffuse attention, a fixed token budget can result in under-selection, as a larger number of tokens is necessary to ensure accurate attention modeling. (b) Current algorithms suffer from varying degrees of inefficiency. As shown in Figure 2, we observe that the saturation point is highly dependent on the specific algorithm, necessitating offline calibration to determine the appropriate budget for each algorithm individually. The actual algorithms, like Quest (Tang etĀ al., 2024b) or DS (Yang etĀ al., 2024c), have to over-select some tokens as the inevitable inaccuracy in estimating the importance of tokens comparing with oracle.
In this work, we reveal that top- methods exhibit issues similar to those previously encountered in LLM sampling. Drawing on this analogy, we introduce top- sampling into sparse attention to address the budget estimation problem. Our study demonstrates that top- can determine the KV cache budget in a more intrinsic and dynamic way compared to top-. Based on these observations, we built Twilight, a hierarchical KV cache pruning framework that enhances existing sparse attention algorithms with adaptive budgeting capabilities. Specifically, Twilight first lets the base algorithm to select a large subset of tokens using a conservative budget, and then further refines this subset by retaining only the top- tokens.
Evaluations are conducted in two aspects: accuracy and efficiency. First, we demonstrate that Twilight optimizes the base algorithms with nearly no accuracy loss on both mid-context benchmarks (GSM8K (Cobbe etĀ al., 2021), COQA (Reddy etĀ al., 2019), PG19 dataset (Rae etĀ al., 2019)) and a comprehensive long-context benchmark, Longbench (Bai etĀ al., 2024). Next, we show that Twilight accelerates full attention by up to times and existing sparse attention methods, by up to times, leading to a end-to-end speedup. Our contributions are summarized as follows:
ā¢
We conduct an in-depth investigation into a significant issue in top- sparse attention: the difficulty in identifying the optimal budget (the saturation point). We propose using top- sampling to dynamically determine this point at runtime.
ā¢
We introduce Twilight, a framework that can endow any existing sparse attention method with adaptive budgeting capabilities, thereby improving their efficiency and facilitating their deployment.
ā¢
We evaluate Twilight in terms of both accuracy and efficiency, demonstrating an speedup on self-attention.
2 Related Works
Sparse Attention. Sparse attention techniques aim to reduce the memory and computational burden of self-attention in long-context LLMs. H2OĀ (Zhang etĀ al., 2023), StreamingLLMĀ (Xiao etĀ al., 2023) and SnapKV Ā (Li etĀ al., 2024) evict non-critical tokens in a static, query- agnostic manner, which are often referred to as KV cache compression. In contract, SparQĀ (Ribar etĀ al., 2023), Double SparsityĀ (Yang etĀ al., 2024c) and QuestĀ (Tang etĀ al., 2024b) retain all tokens in GPU and select critical tokens to save loading. Recent works like RetrievalAttentionĀ (Liu etĀ al., 2024) adopts advanced algorithm to better estimate the token criticality. However, these methods are all top- based, requiring configuring an appropriate budget beforehead, leading to over-selection or under-selection. In this paper, we focus on solving this problem via a more essential approach.
Dynamic Budget. Recent studies have extensively demonstrated that the optimal budgets vary significantly across different layers (Cai etĀ al., 2024; Yang etĀ al., 2024a), attention heads (Feng etĀ al., 2025; Xiao etĀ al., 2024; Tang etĀ al., 2024a), and prompts (tasks) (Zhou etĀ al., 2024). These works tend to focus on only one aspect of the dynamism in attention mechanisms. However, in this paper, we point out that it is the different distributions of attention weights that are the root cause of this dynamism
Other KV Cache Optimizations Several alternative approaches focus on optimizing the KV cache beyond sparsification, including quantizationĀ (Hooper etĀ al., 2024; Zirui Liu etĀ al., 2023; Kang etĀ al., 2024; Nawrot etĀ al., 2024), linear attentionĀ (Wang etĀ al., 2020; Katharopoulos etĀ al., 2020), and memory-efficient attention mechanisms such as FlashAttentionĀ (Dao, 2024) and SageAttentionĀ (Zhang etĀ al., 2025, 2024b). Our approach is orthogonal to these methods, as it focuses on adaptive KV cache pruning rather than compression, approximation, or memory-efficient computation, and can be combined with them for enhanced performance.
3 Bringing Top- Sampling to Sparse Attention
In this section, we formulate the current sparse attention methods and re-examine the root causes of the problems. We argue that to mathematically approximate the attention output, the goal is to select a minimal set of indices such that the sum of their attention scores meets a certain threshold. Therefore, we propose that top- sampling should be used instead of top- to filter out the critical tokens.
3.1 Problem Formulation
We start by formulating the sparse attention. Consider the attention computation during the decoding phase, where we have the query vector , and the key-value cache . Here, denotes the head dimension, and represents the context length.
Definition 3.1 (Sparse Attention).
Let be the set of selected indices, the output of the sparse attention equals to
where .
To minimize the output error , we need to carefully select the subset of tokens that are used in the sparse attention computation. However, directly optimizing this objective function without loading the full KV cache is challenging. Earlier research has shown that the distribution of V is relatively smooth (Zhao etĀ al., 2024), which implies that the bound is relatively tight.
Therefore, the objective becomes minimizing , which means selecting a subset of tokens that maximizes the sum of attention weights. If we fix the number of the subset, i.e. , then we have the oracle top- attention:
Definition 3.2 (Oracle Top- Sparse Attention).
Given the budget ,
The oracle top- attention serves as a theoretical upperbound of current sparse methods.
3.2 Rethink the Problem of Top-
The Achillesā heel of top- attention, as we described earlier, is the dilemma in determining a uniform budget . A larger leads to inefficiency, while a smaller results in accuracy loss. We find that this predicament is quite similar to the one encountered in the sampling phase of large language models (LLMs). During the sampling phase, the model samples the final output token from a predicted probability distribution. Nucleus sampling (Holtzman etĀ al., 2019), or top- sampling, was proposed to address the problem that top- sampling cannot adapt to different next-word distributions.
Motivated by this insight, we examine the distributions of attention weights more closely. FigureĀ 3 displays two different types of attention weight distributions in real LLMs mentioned in FigureĀ 1. In EquationĀ 2, we demonstrated that the output error can be related to the sum of attention weights. It is straightforward to observe that, when comparing a flat distribution to a peaked one, a greater number of tokens must be selected in the flat distribution to reach the same cumulative threshold. Therefore, we argue that the core reason for budget dynamism is the dynamic nature of attention weight distributions at runtime. Drawing inspiration from top- sampling, we introduce top- sparse attention by directly apply threshold to the sum of attention weights.
Definition 3.3 (Oracle Top- Sparse Attention).
Given the threshold ,
Comparing to top-, top- is advantageous because it provides a theoretical upperbound of error in EquationĀ 2 by . Under this circumstance, top- reduces the budget as low as possible, making it both efficient and adaptive to different distributions.
4 Twilight
In the previous section, we demonstrated that top- attention can adaptively control the budget while ensuring that the sum of normalized attention weights meets a certain threshold . Our primary goal is to use top- to endow more existing algorithms with adaptive attention sparsity, rather than simply inventing another sparse attention, which is motivated by two main reasons: On one hand, despite their budget-related challenges, existing sparse algorithms have achieved significant success in current serving systems Ā (Kwon etĀ al., 2023; Zheng etĀ al., 2023), thanks to their effective token selection strategies. These strategies can be readily reused and enhanced with adaptive sparsity. On the other hand, we anticipate that future sparse attention methods may still employ top- selection. By developing a general solution like ours, we aim to automatically equip these future methods with adaptive attention sparsity, thereby improving their efficiency and adaptability without requiring extensive redesign. Consequently, we initially positioned our system, Twilight, as an optimizer for existing algorithms.
However, deploying top- to different existing sparse attention algorithms faces majorly three challenges, both algorithm-wise and system-wise.
(C1) Not all algorithms are suitable for top-. Top- imposes strict constraints on the layout of attention weights. For example, simply replacing top- with top- in Quest (Tang etĀ al., 2024b) would not work, as Quest performs max pooling on weights with a per-page layout (16 tokens per page). Additionally, some other methods (Yang etĀ al., 2024b; Liu etĀ al., 2024) do not use attention weights to select critical tokens at all.
(C2) Itās harder to estimate weights for top- than top-. The precision requirement of top- is higher than that of top-, because the former requires a certain degree of numerical accuracy while the latter only demands ordinality. TableĀ 1 provides a basic comparison of top-, top-, and full attention. The precision requirement of top- attention lies between the other two, which leads us to reconsider the appropriate precision choice for compressing the cache.
(C3) System-level optimizations are needed. Since our work is the first work introduce top- to attention weights, many algorithms need to be efficiently implemented in hardware, including efforts on both efficient parallel algorithm designs and efficient kernel optimizations.
In Section 4.1, we address C1 by proposing a unified pruning framework for sparse attention. In Section 4.2, we mitigate the runtime overhead of by efficient kernel implementations (top-, SpGEMV, Attention) and 4-bit quantization of cache, addressing C2 and C3. Lastly, in Section 4.3, we analyze the overhead of Twilight and discuss some topics.
| Methods | Efficiency | Precision | Output | Need |
| Requirement | Accuracy | Normalize? | ||
| Top- | High | Low | Median | |
| Top- | High | Median | High | |
| Full Attn. | Low | High | High |
4.1 Hierarchical Pruning with Select-then-Prune Architecture
Recall that existing algorithms face the challenge of choosing a proper budget: either over-selection or under-selection. To address this, we propose a two-step process. We first abstract the base algorithm into a black-box Token Selector, which uses some metadata to select a subset of critical tokens. And we allow the Token Selector to a conservative budget (e.g. sparsity), as we have a Pruner after that to further optimize the selected indices by pruning unimportant tokens. This hierarchical Select-then-Prune architecture is illustrated on the left side of FigureĀ 4.
4.2 Efficient Kernel Implementation
4.2.1 Efficient SpGEMV with 4-bit Quantization of Key Cache
As previously analyzed, the precision requirement of top- lies between top- and full attention. For top-, many works (Yang etĀ al., 2024c; Zhang etĀ al., 2024a) push the compression to a extreme low-bit (1-bit/2-bit). For full attention, SageAttention (Zhang etĀ al., 2025) is proposed recently as a 8-bit accurate attention by smoothing and per-block quantization. In this work, we find 4-bit strikes a balance between accuracy and efficiency, making it the ideal choice to calculate estimated attention weights for top-. And we implement an efficient sparse GEMV (SpGEMV) kernel based on FlashInfer. We maintain an extra INT4 asymmetrically quantized cache in GPU as FigureĀ 4 shows. The INT4 vectors are unpacked and dequantized in shared memory which reduces I/O between global memory and shared memory to at most , leading to a considerable end-to-end speedup.
4.2.2 Efficient Top- Kernel via Binary Search
As we mentioned before, our top- method is motivated by the top- sampling, which also takes up a portion of decode latency. Therefore, our efficient kernel is modified from the top- sampling kernel from FlashInfer (Ye etĀ al., 2025), a high performance kernel library for LLM serving.
A brute-force way to do top- sampling is to sort the elements by a descending order and accumulate them until the sum meets the threshold, which is quite inefficient in parallel hardwares like modern GPU. Our kernel adopts a parallel-friendly binary search algorithm as AlgorithmĀ 1.
4.3 Overhead Analysis and Discussion
Runtime Overhead. The runtime of Twilight-optimized algorithm consists of three parts according to the pipeline in FigureĀ 4: . Comparing to the baseline without Twilight, our method introduces an extra latency term but reduces because it further reduces its I/O. Our hierarchical architecture naturally fits the hierarchical sparsity, where the number of tokens gradually decreases as the precision increases. Suppose the Token Selector has a sparsity, then the theoretical speed up can be formulated as
where is the budget of Token Selector, is the budget after pruned by Twilight. Suppose , then the speed up is approximately . Here we omit the overhead of top- since SpGEMV dominates the latency when .
Integrate with Serving System. Since our system design naturally aligns with PagedAttention (Kwon etĀ al., 2023), Twilight can be seamlessly integrated into popular serving systems like vLLMĀ (Kwon etĀ al., 2023) and SGLangĀ (Zheng etĀ al., 2023). Prefix sharing and multi-phase attention (Lin etĀ al., 2024; Zheng etĀ al., 2023; Zhu etĀ al., 2024; Ye etĀ al., 2024a, b) also become common techniques in modern serving systems, which also fit Twilight since we use paged-level or token-level sparse operations and can achieve flexible computation flow.
5 Evaluation
In this section, we will perform several experiments to demonstrate that equipping current SOTA sparse attention algorithms with Twilight improves their efficiency while preserving the accuracy.
5.1 Accuracy Evaluation
5.1.1 Setup
Benchmarks and Models. We evaluate Twilight on two types of benchmarks: Longbench (Bai etĀ al., 2024) for long context and three separate tests for short context (k tokens), GSM8K (Cobbe etĀ al., 2021), COQA (Reddy etĀ al., 2019) and perplexity on PG19 dataset (Rae etĀ al., 2019). We select three widely used models, Longchat-7B-v1.5-32k (Li etĀ al., 2023), LLaMA2-7b-chat (Touvron etĀ al., 2023) and LLaMA-3.1-8B-Instruct (Dubey etĀ al., 2024) (128k context length), with two of them have the long context ability 32k, covering two mainstream attention implementations multi-head attention (MHA) and group query attention (GQA).
Baselines. We use two SOTA top- sparse attention methods Quest (Tang etĀ al., 2024b) and DS (Yang etĀ al., 2024c). The hyperparameter of Twilight is set to for LLaMA 2/3 and for Longchat. We do not apply any sparse methods to the first two layers as it is illustrated in baselinesā paper to ensure a fair comparison.
5.1.2 Results on Longbench
We comprehensively evaluate Twilightās long context ability on 12 different tasks chosen from Longbench, covering all task types, on two long context models. For each baseline, we choose four budget settings: 256, 1024, 4096, 8192 to observe the impact on budget, and a specially setting named "Twilight", means using Twilight pruner in this algorithm to dynamically determine the budget. In this case, the base algorithm uses a very conservative budget, 8192, in our case. We also equip Full with Twilight (i.e. The token selector is a trivial one which selects all tokens each time) for better comparison.
The full results are shown in TableĀ 3 in AppendixĀ A. In Longchat, the Twilight series outperforms its original version by up to 5.7% in score, while successfully pruning up to 98% of the redundant tokens over-selected by the base algorithm. In LLaMA-3.1-8b-Instruct, Twilight achieves almost zero accuracy loss (1%) with a slight increase in budget usage. We hypothesize that this slight increase is due to the knowledge being more compressed in LLaMA-3.1.
5.1.3 Results on Short Tasks
| Methods | GSM8K(flexible/strict) | COQA(em/f1) | PG-19 Perplexity |
| Llama-2-7B-Chat | |||
| Full Cache | 0.2290/0.2282 | 0.5935/0.7511 | 7.503 |
| Quest | 0.0114/0.0061 | 0.5150/0.6991 | 14.15 |
| DS | 0.1820/0.1812 | 0.6043/0.7632 | 7.622 |
| Twilight | 0.2153/0.2115 | 0.6088/0.7642 | 7.600 |
| Llama-3.1-8B-Instruct | |||
| Full Cache | 0.7726/0.7475 | 0.6363/0.7882 | 7.490 |
| Quest | 0.0773/0.0652 | 0.5310/0.7033 | 19.00 |
| DS | 0.3806/0.3609 | 0.6490/0.8003 | 10.23 |
| Twilight | 0.7771/0.7604 | 0.6325/0.7869 | 7.529 |
We then demonstrate that the Twilight pruner itself does not negatively impact performance on two zero-shot generation tasks: GSM8K and COQA using lm-harness framework (Gao etĀ al., 2021), as well as one perplexity test in the PG-19 data set. Since we are specifically evaluating the pruner, we do not integrate Twilight into the baseline models. All baselines use a budget of 64, which is comparable to the budget after Twilightās pruning. The results in TableĀ 2 show that Twilight outperforms Quest and DS by a significant margin, with nearly zero loss compared to full attention.
5.2 Efficiency Evaluation
5.2.1 Setup
Datasets. We evaluate the efficiency of Twilight on both the self-attention operator and the end-to-end decoding latency on a single A100 GPU. Since the budget is prompt (task)-related, our experiments are conducted on Longbench, from which we select three different types of tasks: Qasper (Dasigi etĀ al., 2021) for QA, GovReport (Huang etĀ al., 2021) for summarization, and LCC (Guo etĀ al., 2023) for coding. We use prompts ranging from 10k to 30k tokens for evaluation. Given that Twilight is designed for deploying sparse attention in serving systems, we use batch inference in our experiments.
Baselines and Implementation Details. We compare our methods with the following baselines: PyTorchās scaled-dot-product-attention (SDPA), with FlashAttention2 (Dao, 2024) and Memory-Efficient Attention (Lefaudeux etĀ al., 2022) as backends. FlashInfer (Ye etĀ al., 2025), a high-performance kernel library for LLM serving. Quest, which achieves state-of-the-art latency performance among sparse attention methods. We integrate Twilight with both FlashInfer and Quest, resulting in FlashInfer-Twi and Quest-Twi. We modified the Quest kernels to support batch inference. We implemented Twilight using both CUDA and OpenAI Triton (Tillet etĀ al., 2019), with the technical details described in Section 4.2.
5.2.2 Speedup on Self-Attention
We first evaluate the speedup on self-attention operator across different batch sizes and sequence length. As FigureĀ 7 shows, FlashInfer-Twi and Quest-Twi achieve a speedup up to and separately comparing with FlashAttention2. Moreoever, they accelerates the base algorithm (FlashInfer and Quest) by and , separately.
5.2.3 End-to-End Decoding Speedup
We conduct end-to-end inference evaluation with the similar settings, with the difference that the batch size ranges from 32 to 256, for a better simulation of serving scenario. FigureĀ 8 tells that Quest-Twi achieves up to decoding acceleration comparing with FlashInfer, with a extra speedup comparing to Quest without Twilight.
5.3 Ablation Study
Selection of Quantization Bits. We previously argued that 4-bit quantization is suitable for estimating attention weights in Section 4.2. To empirically verify this, we examined the sum of real attention weights for 2-bit, 4-bit, and 8-bit quantizations. As shown in FigureĀ 5, the sum of attention weights for 2-bit quantization drops significantly, while 4-bit and 8-bit quantizations maintain stability. Although 8-bit quantization can be used for attention computation (Zhang etĀ al., 2025), it is overly precise for the purpose of top- sampling. Considering the balance between estimation accuracy and computational efficiency, 4-bit quantization emerges as a good choice.
Time Breakdown for Twilight. Given Twilightās hierarchical architecture, which comprises three distinct components, it is crucial to analyze the time breakdown. FigureĀ 6 illustrates the time breakdown for different batch sizes in a 32k retrieval task. In this scenario, Quest employs a budget of 8192 (1/4 sparsity), while Twilight further prunes this budget down to 256. The breakdown aligns closely with the theoretical cost model we presented earlier, demonstrating that Twilight significantly reduces the time required for the sparse attention kernel while introducing some additional overheads.
6 Conclusion
In this paper, we first highlight that top- attention struggles to find an optimal budget due to the dynamic nature of attention weight distributions. We then introduce Twilight, a framework with hierarchical selector-optimizer architecture that leverages top- filtering to address this challenge. Empirical results demonstrate that Twilight can adaptively prune up to 98% of redundant tokens, resulting in a acceleration in self-attention operations and a reduction in end-to-end per-token latency. Comparing to the base algorithm it is applied to, Twilight offers an additional speedup. Our work underscores the importance of adaptive attention sparsity, paving the way for future research on efficient sparse attention.
Appendix A Full Results on Longbench
| Methods | Budget | Single-Doc. QA | Multi-Doc. QA | Summarization | Few-shot | Code | Synthetic | Avg. Score | ||||||
| Qasper | MF-en | HotpotQA | 2WikiMQA | Musique | GovReport | QMSum | MultiNews | TriviaQA | LCC | Repobench-P | PR-en | |||
| Longchat-7B-32k | ||||||||||||||
| Full | 32k | 29.48 | 42.11 | 30.97 | 23.74 | 13.11 | 31.03 | 22.77 | 26.09 | 83.25 | 30.50 | 52.70 | 55.62 | 36.78 |
| Twilight (Avg. 146) | 31.74 | 43.91 | 33.59 | 25.65 | 13.93 | 32.19 | 23.15 | 26.30 | 85.14 | 34.50 | 54.98 | 57.12 | 38.52(+4.7%) | |
| Quest | 256 | 26.00 | 32.83 | 23.23 | 22.14 | 7.45 | 22.64 | 20.98 | 25.05 | 67.40 | 33.60 | 48.70 | 45.07 | 31.26 |
| 1024 | 31.63 | 42.36 | 30.47 | 24.42 | 10.11 | 29.94 | 22.70 | 26.39 | 84.21 | 34.5 | 51.52 | 53.95 | 36.85 | |
| 4096 | 29.77 | 42.71 | 32.94 | 23.94 | 13.24 | 31.54 | 22.86 | 26.45 | 84.37 | 31.50 | 53.17 | 55.52 | 37.33 | |
| 8192 | 29.34 | 41.70 | 33.27 | 23.46 | 13.51 | 31.18 | 23.02 | 26.48 | 84.70 | 30.00 | 53.02 | 55.57 | 37.10 | |
| Twilight (Avg. 131) | 31.95 | 43.28 | 31.62 | 24.87 | 13.48 | 32.21 | 22.79 | 26.33 | 84.93 | 33.50 | 54.86 | 56.70 | 38.04(+2.5%) | |
| DS | 256 | 28.28 | 39.78 | 27.10 | 20.75 | 9.34 | 29.68 | 21.79 | 25.69 | 83.97 | 32.00 | 52.01 | 53.44 | 35.32 |
| 1024 | 30.55 | 41.27 | 30.85 | 21.87 | 7.27 | 26.82 | 22.95 | 26.51 | 83.22 | 31.50 | 53.23 | 55.50 | 35.96 | |
| 4096 | 28.95 | 41.90 | 32.52 | 23.65 | 8.07 | 29.68 | 22.75 | 26.55 | 83.34 | 30.00 | 52.77 | 55.48 | 36.31 | |
| 8192 | 29.05 | 41.42 | 31.79 | 22.95 | 12.50 | 30.44 | 22.50 | 26.43 | 83.63 | 30.50 | 52.87 | 55.33 | 36.62 | |
| Twilight (Avg. 126) | 32.34 | 43.89 | 34.67 | 25.43 | 13.84 | 31.88 | 23.01 | 26.32 | 85.29 | 35.50 | 55.03 | 57.27 | 38.71(+5.7%) | |
| Llama-3.1-8B-Instruct | ||||||||||||||
| Full | 128k | 46.17 | 53.33 | 55.36 | 43.95 | 27.08 | 35.01 | 25.24 | 27.37 | 91.18 | 99.50 | 62.17 | 57.76 | 52.01 |
| Twilight (Avg. 478) | 43.08 | 52.99 | 52.22 | 44.83 | 25.79 | 34.21 | 25.47 | 26.98 | 91.85 | 100.00 | 64.06 | 58.22 | 51.64(-0.7%) | |
| Quest | 256 | 24.65 | 37.50 | 30.12 | 23.60 | 12.93 | 27.53 | 20.11 | 26.59 | 65.34 | 95.00 | 49.70 | 45.27 | 38.20 |
| 1024 | 38.47 | 49.32 | 47.43 | 38.48 | 20.59 | 33.71 | 23.67 | 26.60 | 81.94 | 99.50 | 60.78 | 52.96 | 47.79 | |
| 4096 | 43.97 | 53.64 | 51.94 | 42.54 | 24.00 | 34.34 | 24.36 | 26.75 | 90.96 | 99.50 | 62.03 | 55.49 | 50.79 | |
| 8192 | 44.34 | 53.25 | 54.72 | 44.84 | 25.98 | 34.62 | 24.98 | 26.70 | 91.61 | 100.00 | 62.02 | 54.20 | 51.44 | |
| Twilight (Avg. 427) | 43.44 | 53.2 | 53.77 | 43.56 | 25.42 | 34.39 | 25.23 | 26.99 | 91.25 | 100.0 | 63.55 | 58.06 | 51.57(+0.3%) | |
| DS | 256 | 38.24 | 49.58 | 43.38 | 31.98 | 15.52 | 33.40 | 24.06 | 26.86 | 84.41 | 99.50 | 53.28 | 48.64 | 45.74 |
| 1024 | 42.97 | 54.65 | 51.75 | 33.92 | 20.39 | 34.50 | 24.92 | 26.71 | 92.81 | 99.50 | 62.66 | 48.37 | 49.43 | |
| 4096 | 43.50 | 53.17 | 54.21 | 44.70 | 23.14 | 34.73 | 25.40 | 26.71 | 92.78 | 99.50 | 62.59 | 51.31 | 50.98 | |
| 8192 | 43.82 | 53.71 | 54.19 | 45.13 | 23.72 | 34.27 | 24.98 | 26.69 | 91.61 | 100.00 | 62.40 | 52.87 | 51.14 | |
| Twilight (Avg. 446) | 43.08 | 52.89 | 54.68 | 44.86 | 24.88 | 34.09 | 25.20 | 27.00 | 91.20 | 100.00 | 63.95 | 58.93 | 51.73(+1.2%) | |